# M3RS: Multi-robot, Multi-objective, and Multi-mode Routing and Scheduling

### v2

@@ IV-B3 Results
The mean and standard deviation of the performance metrics for solutions of RS-F and M3RS variants are presented in Table III. RS-F (Min) achieves the highest success rate ($SR$), addressing all tasks. This is because the lowest disinfection dosage is used by default, reducing disinfection time and increasing $SR$, but at the cost of lower disinfection quality ($DQ$). Conversely, RS-F (Max) provides higher $DQ$ but lower $SR$. The higher disinfection quality requires more time, limiting agents’ ability to complete other tasks. The solutions of M3RS outperform the RS-F baselines in the combined metric ($MSI$), offering a better balance between $SR$ and $DQ$. The solutions of M3RS admit better $DQ$ and $ATQ$ scores. The $QBR$ scores show that solutions of M3RS offer better quality improvement relative to the additional resource expenditure (energy and time). This suggests that considering disinfection quality through variable disinfection modes improves overall disinfection quality while maintaining or improving task completion rates.
@@
We further analyze the missions by looking at the ratio of idle time to task servicing time, as well as the $SR$ and $DQ$ for different time-window types (see Fig. 2 (a)-(c)). The solution of M3RS show less idle time and more working time, indicating better use of mission time, see Fig. 2 (a)-(b) . Using variable disinfection modes likely saves time and energy, which can be spent on disinfecting more tasks or improving quality for others, depending on the user preference i.e. $\lambda$. RS-F (Max) shows higher idle time and lower working time than M3RS, as disinfection in the highest modes requires robots to arrive at or before a task’s time window. Early arrivals lead to waiting, limiting task completion and resulting in lower SR and DQ compared to M3RS.
@@
When examining $SR$ and $DQ$ across time-window types, tasks available throughout the mission have the best performance, while those with limited time windows perform worse, see Fig. 2 (c). This suggests that relaxing time constraints where possible could lead to better overall performance.
@@ IV-C1 Baselines
The performance of solving M3RS with CCG is compared with that of directly solving using MIP solver. For CCG, we report results with varying number of clusters i.e $K\in[2,4,6,8]$. The performance is evaluated in terms of the relative optimality gap ($OG$) given by Eq (30). For CCG, we use the best bound from the MIP solver to compute the $OG$.
@@ IV-C2 Setup
For this, four problem types are considered with a number of tasks and robots mentioned in Table. IV. For each problem type, 5 instances are generated using the procedure outlined in Sec. IV-B. User preference for each instance is sampled from $\lambda\in[0.1,0.9]$, with higher $\lambda$ values prioritizing task completion over disinfection quality. We use load factors $\eta=1$ and $\eta=2$, where higher $\eta$ simulates missions with shorter $T_{H}$, providing less time for mission completion. The mission time ($T_{H}$) is calculated using Eq (24). The time limit for MIP solver and CCG is set to 10 minutes. We further report the performance with time limit for MIP set to 25 minutes.
@@ IV-C3 Discussion
The mean and standard deviation of the $OG$ metric for MIP and CCG with time limit set to ten minutes is reported in Table IV. The change in $OG$ for MIP solver relative to CCG, with time limit for MIP solver set to 25 minutes is reported in Table V. The $OG$ achieved by CCG solutions is consistently better than that of MIP solver with the same time limit, see Table IV. Further, it is seen with increased time limit the performance of CCG remains comparable or better than MIP solver while being 2.5×2.5$\times$2.5 × faster. This indicates that clustering and sampling approach coupled with column generation provides an effective way to find good solution in shorter time as compared to MIP solver. Generally, clustering improves $OG$ values, with more clusters benefiting larger instances.
Table V: Difference of average optimality gap ($OG$) between CCG and the MIP solver across various instances. A 25-minute time limit is set for the MIP solver. Positive differences indicate better $OG$ for the MIP solver, while negative values favor CCG. Colors represent MIP outperforming CCG and CCG outperforming MIP. CCG achieves comparable or better $OG$ while being approximately 2.5×2.5$\times$2.5 × faster than the MIP baseline.
Metric $\eta$ $\lambda$ $120-4$ $150-5$ $175-5$ $200-6$
$OG_{CCG}-OG_{MIP}$ $1.0$ $0.9$ $0.038$ $-0.084$ $-0.058$ $-0.077$
$0.1$ $0.022$ $-0.052$ $-0.007$ $-0.026$
$2.0$ $0.9$ $0.003$ $-0.154$ $-0.074$ $-0.150$
$0.1$ $-0.029$ $-0.080$ $-0.091$ $-0.105$
@@ IV-D Simulated and Hardware Case Study
(c) The route sequences for simulated mission with 20 tasks, 2 robots, $T_{H}$ of 0.4 hours and $\lambda=0.9$.
(d) The route sequence for hardware experiment with $\lambda=0.9$
(e) The route sequence for hardware experiment with $\lambda=0.1$
Figure 3: (a) A team of G-robots [32] disinfecting tables. (b) The routes of two robots executing various tasks in the simulated environment. The route of the first robot is highlighted in white and for the second robot yellow is used. (c) The route sequence taken by each robot in the simulated mission. For $\lambda=0.9$, maximization of the number of tasks completed has more emphasis than the quality of disinfection. As a result, most of the tasks are done in higher modes. (d) & (e) The route sequence taken by each robot for the hardware mission with $\lambda=0.9$ and $0.1$. For $\lambda=0.9$, the emphasis is on disinfecting all the tables at higher disinfection modes. Whereas, for $\lambda=0.1$, the emphasis is on ensuring high-quality disinfection rather than disinfecting all tables.
IV-D1 Simulation and Hardware Experiment Setup
Multiple disinfection mission instances are generated for a simulated environment. Each disinfection mission is characterized by mission time, time windows for each task, and number of robots used in the mission. The simulated environment consists of $20$ tables as disinfection surfaces spread across two rooms. The fleet size considered in the range, $|\mathcal{A}|=2-5$. The mission time is calculated using Eq. 24. The entire simulated experiment is conducted in ROS [35] and Gazebo environment [36]. A simulated mission is depicted in Fig 3 (b).
For the hardware case study, we use two G-robots (shown in Fig. 3 (a)) for a disinfection mission consisting of six tables. Each table can be disinfected in three disinfection modes, i.e. $D_{99.99}$, $D_{99}$, and $D_{90}$. The entire disinfection mission must be completed in $17$ minutes.
For both experiments, M3RS is solved for two preferences: $\lambda\,=0.9$ and $0.1$ for each mission. The optimization model is solved in CPLEX [33]. The metrics reported include $SR$, $DQ$, $MSI$, and average lateness of the robot in task completion.
@@ IV-D2 Discussion
The metrics for simulated and hardware case studies are reported in Table VI. The example of generated route sequences in simulated experiment is visualized in Fig. 3 (b)-(c). The route sequence for hardware experiments is shown in Fig. 3 (d)-(e). From Table VI, it can be seen that a decent $SR$ is achieved in all missions. Even for missions with tight $T_{H}$ the $SR$ is high. Therefore, a larger robotic fleet would be necessary to compensate for tighter mission times in order to service as many tasks as possible. For missions with shorter $T_{H}$, there is a decline in $DQ$ metric. This is because with tighter mission times, each task has a smaller time window, which means many of the tasks cannot be performed in higher disinfection modes. A possible remedy could be to enable robots to collaborate on different tasks. Finally, there are some minor delays observed in terms of task completion times for simulation and hardware implementation. These delays are caused by slightly longer path navigation times. Such delays can be accounted for by introducing chance constraints to model stochasticity in the system.
Table VI: Results for simulated and hardware experiments.
Num. of tasks and robots $T_{H}$ (hrs) $\lambda$ $SR$ $DQ$ MSI Avg. Lateness (s)
@@ V Conclusions and Future work
In this work, we proposed M3RS, a novel multi-objective formulation that routes and schedules a team of agents to accomplish various tasks in a mission. The agents can perform the task in one of many available modes. Each mode has distinct task quality, resource consumption, and service times. We first propose a mixed-integer linear programming model for M3RS. Then, a column generation approach is presented to tackle larger instances with a faster computation time. We showcase the utility of M3RS in multi-robot disinfection applications. The experiments suggest that our solution for M3RS provides more flexibility to users to select solutions based on desired preferences. The current approach is limited as it assumes a deterministic environment, unlike real-world scenarios. Generating the entire Pareto front is computationally expensive, particularly for larger instances. Our future work will extend M3RS to incorporate elements like stochastic travel times, dynamic arrival of tasks and realistic issues like robot breakdown. Further, we aim to tightly integrate the robot path planning with the decision making framework provided by M3RS. Finally, we will investigate using deep reinforcement learning techniques to generate good quality solutions in real-time.
Acknowledgements

### v3

@@ IV-B3 Results
Fig. 4 presents the box-plots of performance metrics for RS-F and M3RS variants. RS-F (Min) uses the lowest disinfection mode. This minimizes execution time and maximizes success rate ($SR$), due to reduced task completion time. However, it results in the lowest disinfection quality ($DQ$). RS-F (Max) defaults to the highest disinfection mode. This increases execution time and prioritizes disinfection quality. However, it lowers the success rate due to longer disinfection times. M3RS ($\lambda=0.01$) improves disinfection quality by 10% over RS-F (Max) and 38% over RS-F (Min) while maintaining a strong balance between efficiency and quality. In terms of the combined metric ($MSI$), M3RS ($\lambda=0.01$) achieves 10% and 15% gains over RS-F (Max) and RS-F (Min), respectively, demonstrating its ability to optimize both success rate and disinfection quality. M3RS solutions also achieve higher Average Task Quality ($ATQ$), with a 3% gain over RS-F (Max) and 46% over RS-F (Min). The Quality-Benefit Ratio ($QBR$) improves by 6% compared to RS-F (Max), indicating that M3RS provides better disinfection quality relative to resource expenditure (energy and time). These results indicate that incorporating variable disinfection modes yields better overall disinfection quality while maintaining or improving task completion rates. To further ascertain the gains made by task execution modes, we conducted a statistical comparison of success rate $SR$ and disinfection quality $DQ$ between RS-F (Max) and M3RS ($\lambda=0.01$) using Welch’s t-test (Sakai and Sakai,, 2018). This test was applied to evaluate whether the differences observed in each metric between the two variants are statistically significant. For the $SR$, the p-value was $7.74\times 10^{-6}$, and for the $DQ$, the p-value was $5.08\times 10^{-6}$. Both results indicate statistically significant differences at the $95\%$ confidence level, demonstrating that M3RS ($\lambda=0.01$) achieves significant improvements over RS-F (Max) in these metrics.
(a) Box-plot comparing idling time to $T_{H}$ ratio for M3RS and RS-F for different number of tasks and a fleet size of 4.
(b) Box-plot comparing working time to $T_{H}$ ratio for M3RS and RS-F for different number of tasks and a fleet size of 4.
(c) Box-plot depicting impact on SRSR and $DQ$ for different types of time windows. The tasks with time-windows could be critical or non-critical.
Figure 5: (a) & (b) The solution of M3RS makes effective use of agents as indicated by lower idling times and higher working times. (c) Further, tasks available during the entire mission have a higher task completion rate and disinfection quality.
@@
We further analyze the missions by looking at the ratio of idle time to task servicing time, as well as the $SR$ and $DQ$ for different time-window types (see Fig. 5 (a)-(c)). The solution of M3RS shows less idle time and more working time, indicating better use of mission time; see Fig. 5 (a)-(b). Use of variable disinfection modes promotes better use of time and energy, which could be spent on disinfecting more tasks or improving quality for others, depending on the user preference i.e. $\lambda$. In most cases, RS-F (Max) shows higher idle time and lower working time than M3RS, as disinfection in the highest modes requires robots to arrive at or before a task’s time window. Early arrivals lead to waiting, which limits task completion and results in lower SR and DQ compared to M3RS. For missions with more than 50 tasks, we observe that for certain $\lambda$ values, M3RS shows slightly higher average idling times and lower average working times compared to RS-F (Max). This could be attributed to the increased problem size with more tasks and may be improved with more efficient solution strategies.
@@
When examining $SR$ and $DQ$ across time-window types, tasks available throughout the mission have the best performance, while those with limited time windows perform worse, see Fig. 5 (c). This suggests that relaxing time constraints, where possible, could lead to better overall performance.
Overall, the synthetic experiments demonstrate that incorporating quality as a decision variable in M3RS enables more effective planning of disinfection missions, balancing the number of tasks completed with the quality of service within given time and resource constraints. While this work focuses on disinfection, M3RS is also applicable to other domains, such as cleaning robots, where balancing cleaning efficiency and time is critical (Megalingam et al.,, 2025), and robot patrolling, which involves managing trade-offs between high- and low-priority areas (Katole et al.,, 2023).
@@ IV-C Performance of CCG
We evaluate our clustering-based column generation (CCG) method for solving M3RS by comparing it with a standard MIP solver (Laborie et al.,, 2018). The results show that CCG achieves comparable or better performance in less time.
@@ IV-C1 Baselines
The performance of solving M3RS with CCG is compared with that of directly solving using an MIP solver. For CCG, we report results with varying number of clusters i.e $K\in[2,4,6,8]$. The performance is evaluated in terms of the relative optimality gap ($OG$) given by Eq. (32). For CCG, we use the best bound from the MIP solver to compute the $OG$.
@@ IV-C2 Setup
For this, four problem types are considered with a number of tasks and robots mentioned in Fig. 6. For each problem type, 10 instances are randomly generated using the procedure outlined in Sec. IV-B. User preference for each instance is sampled from $\lambda\in[0.1,0.9]$, with higher $\lambda$ values prioritizing task completion over disinfection quality. We use load factors $\eta=1$ and $\eta=2$, where a higher $\eta$ simulates missions with shorter $T_{H}$, providing less time for mission completion. The mission time ($T_{H}$) is calculated using Eq. (26). The time limit for the MIP solver and CCG is set to 10 minutes. We further report the performance with a time limit for MIP set to 25 minutes.
(a) $\eta=1.0$ and $\lambda=0.9$
(b) $\eta=2.0$ and $\lambda=0.9$
(c) $\eta=1.0$ and $\lambda=0.1$
(d) $\eta=2.0$ and $\lambda=0.1$
Figure 6: (a)-(d) Box-plots comparing the optimality gap ($OG$) of M3RS problems solved by: (i) a standard MIP solver (Laborie et al.,, 2018), and (ii) our Clustering-based Column Generation (CCG) approach. CCG is evaluated with different numbers of clusters ($K=2,4,6,8$). Both methods are given a 10-minute time limit. Box-plots are generated from multiple randomly generated problem instances, varying in $\lambda$ and $\eta$. Lower $OG$ values indicate better performance. CCG consistently achieves a lower OGOG than the MIP solver within the same time budget.
Table III: Average optimality gap ($OG$) difference between CCG and the MIP solver across various instances, with a 25-minute time limit for MIP and a 10-minute time limit for CCG. Positive values indicate better OGOG for MIP, while negative values favor CCG. All values are given in $\%$. Colors highlight MIP outperforming CCG and CCG outperforming MIP. Our CCG method achieves comparable or better OGOG while using 60% less computation time than the MIP solver.
Metric $\eta$ $\lambda$ # tasks - # robots.
1204120-4 1505150-5 1755175-5 2006200-6
$OG_{CCG}-OG_{MIP}$ $1.0$ $0.9$ $4.1\%$ $7.2\%$ $3.2\%$ $6.0\%$
$0.1$ $5.0\%$ $3.0\%$ $0.9\%$ $0.6\%$
$2.0$ $0.9$ $0.7\%$ $15.7\%$ $4.2\%$ $15.4\%$
$0.1$ $0.1\%$ $8.6\%$ $6.2\%$ $9.2\%$
@@ IV-C3 Discussion
[…]
@@ IV-D Simulated and Hardware Case Study
In this section, simulation and hardware case studies are presented which apply M3RS in realistic scenarios. Videos of the experiments are available on the project page: https://sites.google.com/view/g-robot/m3rs/.
IV-D1 Simulation and Hardware Experiment Setup
Multiple disinfection mission instances are generated for a simulated environment. Each disinfection mission is characterized by mission time, time windows for each task, and the number of robots used in the mission. The simulated environment consists of $20$ tables as disinfection surfaces spread across two rooms. The fleet size considered in the range, $|\mathcal{A}|=2$ to $5$. The mission time is calculated using Eq. (26). The entire simulated experiment is conducted in ROS (Quigley et al.,, 2009) and the Gazebo environment (Koenig and Hsu,, 2013). A simulated mission is depicted in Fig 7 (b).
For the hardware case study, we use two G-robots (shown in Fig. 7 (a)) for a disinfection mission consisting of six tables. Each table can be disinfected in three disinfection modes, i.e., $D_{99.99}$, $D_{99}$, and $D_{90}$. The entire disinfection mission must be completed in $17$ minutes. For both experiments, M3RS is solved for two preferences: $\lambda\,=0.9$ and $0.1$ for each mission. The optimization model is solved in CPLEX (Laborie et al.,, 2018). The reported metrics are $SR$, DQDQ, MSIMSI, and the average lateness of the robot in task completion.
@@ IV-D1 Simulation and Hardware Experiment Setup
(c) The route sequence for hardware experiment with $\lambda=0.9$
(d) The route sequence for hardware experiment with $\lambda=0.1$
(e) The route sequences for simulated mission with 20 tasks, 2 robots, $T_{H}$ of 0.4 hours and $\lambda=0.9$.
Figure 7: (a) A team of G-robots (Mehta et al.,, 2022) disinfecting tables. (b) The routes of two robots executing various tasks in the simulated environment. The route of the first robot is highlighted in white and for the second robot yellow is used. (c) & (d) The route sequence taken by each robot for the hardware mission with $\lambda=0.9$ and $0.1$. For $\lambda=0.9$, the emphasis is on disinfecting all the tables at higher disinfection modes. Whereas, for $\lambda=0.1$, the emphasis is on ensuring high-quality disinfection rather than disinfecting all tables. (e) The route sequence taken by each robot in the simulated mission. For $\lambda=0.9$, maximization of the number of tasks completed has more emphasis than the quality of disinfection. As a result, most of the tasks are done in higher modes.
@@ IV-D2 Discussion
The metrics for simulated and hardware case studies are reported in Table IV. The example of generated route sequences in a simulated experiment is visualized in Fig. 7 (b) & (e), respectively. The route sequence for hardware experiments is shown in Fig. 7 (c)-(d). From Table IV, it can be seen that a decent $SR$ is achieved in all missions. Even for missions with tight $T_{H}$, the $SR$ is high. Therefore, a larger robotic fleet would be necessary to compensate for tighter mission times to service as many tasks as possible. For missions with shorter $T_{H}$, there is a decline in the $DQ$ metric. This is because with tighter mission times, each task has a smaller time window, which means many of the tasks cannot be performed in higher disinfection modes. A possible remedy could be to enable robots to collaborate on different tasks. Finally, delays in task completion times were observed in both simulation and hardware implementations. These are primarily due to longer-than-expected navigation times caused by environmental and planning uncertainties. Such delays can be mitigated by incorporating chance constraints to model system stochasticity more effectively.
Table IV: Results for simulated and hardware experiments.
# tasks - # robots $T_{H}$ (hrs) $\lambda$ SRSR DQDQ MSI Avg. Lateness (s)
@@ V Conclusions and Future work
In this work, we proposed M3RS, a novel multi-objective formulation that routes and schedules a team of agents to accomplish various tasks in a mission. The agents can perform the task in one of many available modes. Each mode has distinct task quality, resource consumption, and service times. We first propose a mixed-integer linear programming model for M3RS. We showcase the utility of M3RS in multi-robot disinfection applications. The experiments suggest that our M3RS demonstrates 3%–46% performance improvements over standard task allocation methods across various metrics. The solutions for M3RS provide more flexibility to users to select solutions based on desired preferences. We also presented a clustering-based column generation approach that solves M3RS instances in $60\%$ lower compute time while maintaining competitive solutions.
The current approach assumes a deterministic environment and simplified robot energy models. Our future work will extend M3RS to incorporate elements like stochastic travel times, dynamic arrival of tasks, realistic robot energy models, and realistic issues like robot breakdown. Further, we aim to tightly integrate the robot path planning with the decision-making framework provided by M3RS. Finally, we will investigate using deep reinforcement learning techniques to generate good-quality solutions in real-time.
Acknowledgments